# The High W Challenge: Robust Neutrino Energy Estimators for LArTPCs

### v1

@@ VII.1 Estimator Comparison in Parameter Exclusion
We compare the performance of different energy estimators in excluding benchmark oscillation hypotheses when specific alternate parameters are assumed. We consider two tests: excluding the global central value of $\Delta m^{2}_{23}=2.451\times 10^{-3}$ eV2 [ParticleDataGroup:2024cfk] using $\nu_{\mu}$ disappearance, and excluding $\delta_{CP}=0$ with $\nu_{e}$ appearance, both assuming DUNE’s far detector flux and baseline.
For each estimator/generator combination, we compute the estimated energy spectra from $\nu_{\mu}$ assuming the central value and the central value $\pm 3\sigma$, where $\sigma=0.026\times 10^{-3}$ eV2 [ParticleDataGroup:2024cfk]. All distributions are normalized to the same arbitrary detector exposure. We then compute the following metric to quantify the strength of the exclusion of the central value:
$\sqrt{X}=\sqrt{\sum_{i}\left(\frac{N^{\prime}_{i}-N_{i}}{\sqrt{N^{\prime}_{i}}}\right)^{2}},$ (13)
where $N_{i}$ is the predicted rate in the $i$th estimated energy bin assuming the central value, and $N^{\prime}_{i}$ is the rate assuming the central value $\pm 3\sigma$. Similarly, in the $\nu_{e}$ appearance case, we treat $\delta_{CP}=0$ as the central value and compute $N^{\prime}_{i}$ using $\delta_{CP}=\pm\frac{\pi}{2}$.
Figure 11 compares the performance of the different estimators relative to a benchmark case: GENIE using the $W^{2}$ estimator. The performance metric is expressed as the ratio of the $\sqrt{X}$ values obtained for each estimator-generator combination to the benchmark case, which allows for a formulation independent of the overall exposure.
For a fixed estimator, there is a substantial spread in performance across generators. By exhibiting a smaller generator-to-generator spread, the SF method appears marginally more robust against model dependence – a result consistent with the smaller spread in bias observed in Fig 5. However, the SF method consistently provides the weakest separation power. For a fixed generator the relative ordering of estimators is fairly consistent, with the calorimetric and $W^{2}$ estimators providing the best separation power.
In $\nu_{\mu}$ disappearance, exclusive methods such as SF or proton-based offer competitive sensitivity, but in $\nu_{e}$ appearance the loss of statistics outweighs gains in robustness. This highlights the importance of inclusive estimators for CP-violation searches.
@@ VII.2 Final State Interactions
where $R(E_{\rm est},E_{\rm true})$ is the response matrix, $P(E_{\rm true};\alpha)$ the oscillation probability, and $\Phi(E_{\rm true})$ the incident flux. The parameter set $\alpha$ denotes the oscillation parameter of interest, i.e. $\delta_{CP}$ in the $\nu_{e}$ appearance channel and $\Delta m^{2}_{32}$ in the $\nu_{\mu}$ disappearance channel. When analyzing the FSI impact, we will only consider $\Delta m^{2}_{32}$ and $\nu_{\mu}$ disappearance; we will explore the impact of mismodelling secondary variables on $\delta_{CP}$ measurements in Section VIII. We inject different assumptions (e.g. FSI on or off) in the estimated energy spectra by replacing the response matrix in equation 14.
@@
$X(\alpha^{\prime})=\sum_{i}\left(\frac{N^{\prime}_{i}(\alpha^{\prime})-N_{i}(\alpha)}{\sqrt{N_{i}(\alpha)}}\right)^{2},$ (15)
@@
Fig. 12 shows the ratio between the extracted value of $\Delta m^{2}_{23}$ obtained from fitting the non-FSI spectrum with equation 15 and the value of $\Delta m^{2}_{23}$ used to generate the full spectrum with FSI. Consistently with Section V.3, the impact of FSI on the oscillation parameters is smallest when using the $W^{2}$ and SF estimators, whereas the use of calorimetric and proton-based methods display sizable biases. The behavior is consistent across GENIE and NuWro.
@@
Figure 12: Bias in extracted $\Delta m^{2}_{32}$ when fitting spectra generated with and without FSI.
@@ VIII Impact of secondary variables dependency
We perform a variation of the closure test described in Section VII.2 to explore the impact of secondary variables (defined in Section VI) on the oscillation parameter extraction. Here, instead of switching off final state interactions to yield a modified response matrix, we construct a series of alternative response matrices by changing the event selection on one of the three secondary variables. We then investigate how the mismodelling of these variables affects the inferred oscillation parameters by fitting the alternative spectrum obtained from the modified response matrix in Eq. 14, and re-minimizing the function in Eq. 15 to extract the corresponding parameter $\alpha^{\prime}$.
@@
We find it is also informative to calculate the shift in the overall bias in neutrino energy across the entire MC simulation sample, $B^{\prime}-B$, as a function of the selection on the secondary variable. These calculations are shown in Figures 17,  15, and 13 and often mimic the bias in the extracted oscillation parameters resulting from the cut on the hidden variable (up to the sign of the bias, as this depends on the oscillation parameter in question).
@@ VIII.1 Hadronic Invariant Mass
Figure 13: Dependence of energy bias on visible hadronic invariant mass. Sharp structures occur near integer multiples of the nucleon mass, reflecting sensitivity to resonance and multi-nucleon production.
@@
(d) $\Delta m^{2}\_{32}$.
Figure 14: Oscillation fits when events are selected by visible hadronic invariant mass. Periodic structures in the extracted $\delta_{CP}$ and $\Delta m^{2}_{32}$ reflect the resonance-like features in Fig. 14.
@@ VIII.2 Missing Hadronic Energy
We explore the impact of missing hadronic energy through constructing the alternative response matrix with an upper limit on this quantity, starting with only events in which all hadronic energy is visible. The results of this closure test are presented in Fig. 16. As expected, the SF method demonstrates the greatest resilience, as this method is able to infer missing momentum from invisible particles, followed by the $W$-based and proton-based estimators. The calorimetric and CCQE-like methods are more sensitive to unobserved hadronic components.
Figure 15: Change in reconstructed energy bias as a function of missing hadronic energy. Methods that incorporate transverse kinematic balance (SF, $W$-based) remain stable, whereas calorimetric and muon-kinematic methods degrade more strongly.
@@
(d) $\Delta m^{2}\_{32}$.
Figure 16: Oscillation fits under restrictions on missing hadronic energy. The SF method yields the most consistent $\delta_{CP}$ and $\Delta m^{2}_{32}$ extraction, though with reduced statistics.
@@ VIII.3 Angular Dependence
We also construct the alternative response matrix by cutting on the lepton scattering angle, starting with only including forward going leptons before loosening the cut to include larger angles, corresponding to more inelastic interactions. Fits of the oscillated spectra are shown in Fig. 18, revealing that the $W$-based estimator is comparatively stable, while the purely CCQE-like approach is highly susceptible to angle mismodelling. The SF method offers the second best performance, likely because this method is applied to selected events that still resemble quasielastic interactions even at large scattering angles.
Figure 17: Change in average bias as a function of a cut on lepton scattering angle, comparing all estimators. Forward-going selections reduce the impact of angular mismodelling, with the $W$-based method showing the most stable behavior.
@@
(d) $\Delta m^{2}\_{32}$.
Figure 18: Extracted oscillation parameters when restricting events by lepton scattering angle. The $W$-based estimator converges rapidly to the correct $\delta_{CP}$ and $\Delta m^{2}_{32}$ values, while the muon-only approach shows strong dependence on $\theta$.
@@ IX Summary
The $W^{2}$ method consistently exhibits the smallest overall bias as functions of the true energy, with a variance comparable to other inclusive estimators, particularly once resolution effects are considered. Additionally, the $W^{2}$ method shows the weakest dependence on FSI effects, most notably when using the GENIE neutrino event generator. When probing the stability of each estimator against the visible hadronic invariant mass, missing hadronic energy, and lepton scattering angle, we find that the $W^{2}$ method often shows the least variation in bias with respect to these variables, especially in the case of the lepton’s scattering angle.
@@
In this work, we examine each method individually, highlighting the contexts and underlying motivations for which each estimator performs best. This thorough examination provides a detailed guidance for future experiments on how to combine estimators to enhance overall performance.

### v2

@@ VII.1 Estimator Comparison in Parameter Exclusion
We compare the performance of different energy estimators in excluding benchmark oscillation hypotheses when specific alternate parameters are assumed. We consider two tests: excluding the global central value of $\Delta m^{2}_{23}=2.451\times 10^{-3}$ eV2 [38] using $\nu_{\mu}$ disappearance, and excluding $\delta_{CP}=0$ with $\nu_{e}$ appearance, both assuming DUNE’s far detector flux and baseline.
For each estimator/generator combination, we compute the estimated energy spectra from $\nu_{\mu}$ assuming the central value and the central value $\pm 3\sigma$, where $\sigma=0.026\times 10^{-3}$ eV2 [38]. All distributions are normalized to the same arbitrary detector exposure. We then compute the following metric to quantify the strength of the exclusion of the central value:
$\sqrt{\chi^{2}}=\sqrt{\sum_{i}\left(\frac{N^{\prime}_{i}-N_{i}}{\sqrt{N^{\prime}_{i}}}\right)^{2}},$ (14)
where $N_{i}$ is the predicted rate in the $i$th estimated energy bin assuming the central value, and $N^{\prime}_{i}$ is the rate assuming the central value $\pm 3\sigma$. Similarly, in the $\nu_{e}$ appearance case, we treat $\delta_{CP}=0$ as the central value and compute $N^{\prime}_{i}$ using $\delta_{CP}=\pm\frac{\pi}{2}$. We compute exclusions varying the oscillation parameters in both directions, though the qualitative conclusions are largely the same.
Figure 12 compares the performance of the different estimators relative to a benchmark case: GENIE using the $W^{2}$ estimator. The ratio of the $\sqrt{\chi^{2}}$ values obtained for two estimator-generator combinations is independent of the overall exposure the distributions are normalized to.
For a fixed estimator, there is a substantial spread in performance across generators. For a fixed generator, the relative ordering of estimators is fairly consistent, with the calorimetric and $W^{2}$ estimators providing the best separation power when examining the $\nu_{\mu}$ spectrum.
We find in both tests the more exclusive methods offer lower sensitivity due to their loss in statistics, with the single proton version of the CCQE-like estimator consistently offering the worse performance. The default CCQE-like and calorimetric estimators perform the best when $\nu_{e}$ appearance is measured, the benefit of their inclusive approach across all energies, followed by the $W^{2}$ method which is the slightly less inclusive given the requirement there is a single proton in the final state. The performance of the default CCQE-like estimator changes substantially when switching from $\nu_{\mu}$ to $\nu_{e}$ spectra, examination of the estimated energy spectra with this estimator shows this is due to worse resolution in the regions sensitive to $\Delta m^{2}_{23}$. The spectra employed to calculate these exclusions along with the contributions to the $\chi^{2}$ statistic bin by bin are shown in the Supplemental Material [1].
@@ VII.2 Final State Interactions
where $R(E_{\rm est},E_{\rm true})$ is the response matrix, $P(E_{\rm true};\alpha)$ the oscillation probability, and $\Phi(E_{\rm true})$ the incident flux. The parameter set $\alpha$ denotes the oscillation parameter of interest, i.e. $\delta_{CP}$ in the $\nu_{e}$ appearance channel and $\Delta m^{2}_{23}$ in the $\nu_{\mu}$ disappearance channel. When analyzing the FSI impact, we will only consider $\Delta m^{2}_{23}$ and $\nu_{\mu}$ disappearance  [^2] The effect of disabling FSI often has a large effect on the shape of the spectrum and this change cannot be offset by adjusting the $\delta_{CP}$ parameter, thus we do not perform this study on $\nu_{e}$ appearance.; we will explore the impact of mismodelling secondary variables on $\delta_{CP}$ measurements in Section VIII. We inject different assumptions (e.g. FSI on or off) in the estimated energy spectra by replacing the response matrix in equation 15.
@@
$\chi^{2}(\alpha^{\prime})=\sum_{i}\left(\frac{N^{\prime}_{i}(\alpha^{\prime})-N_{i}(\alpha)}{\sqrt{N_{i}(\alpha)}}\right)^{2},$ (16)
@@
Fig. 13 shows the ratio between the extracted value of $\Delta m^{2}_{23}$ obtained from fitting the non-FSI spectrum with equation 16 and the value of $\Delta m^{2}_{23}$ used to generate the full spectrum with FSI. The estimators least sensitive to FSI in this test are the two modified versions of the CCQE-like estimator, consistent with the study on bias presented in Section V.3, followed by the $W^{2}$ and SF methods. The calorimetric and proton based methods show the largest biases. The conclusions of this test are consistent across GENIE and NuWro.
@@
Figure 13: Bias in extracted $\Delta m^{2}_{23}$ when fitting spectra generated with and without FSI.
@@ VIII Impact of secondary variables dependency
We perform a variation of the closure test described in Section VII.2 to explore the impact of secondary variables (defined in Section VI) on the oscillation parameter extraction. Here, instead of switching off final state interactions to yield a modified response matrix, we construct a series of alternative response matrices by changing the event selection on one of the secondary variables. We then investigate how the mismodelling of these variables affects the inferred oscillation parameters by fitting the alternative spectrum obtained from the modified response matrix in Eq. 15, and re-minimizing the function in Eq. 16 to extract the corresponding parameter $\alpha^{\prime}$. We do not perform this test with the lepton momentum as this variable strongly correlates with neutrino energy and this test often fails to close as a result.
@@
We find it is also informative to calculate the shift in the overall bias in neutrino energy across the entire MC simulation sample, $B^{\prime}-B$, as a function of the selection on the secondary variable. These calculations are shown in Figures 14,  16, and  18, and often mimic the bias in the extracted oscillation parameters resulting from the cut on the hidden variable (up to the sign of the bias, as this depends on the oscillation parameter in question).
@@ VIII.1 Hadronic Invariant Mass
Figure 14: Dependence the overall bias in neutrino energy as a function of the cut visible hadronic invariant mass. Sharp structures occur near integer multiples of the nucleon mass, reflecting sensitivity to resonance and multi-nucleon production.
@@
(d) $\Delta m^{2}_{23}$.
Figure 15: Oscillation fits when events are selected by visible hadronic invariant mass. Periodic structures in the extracted $\delta_{CP}$ and $\Delta m^{2}_{23}$ reflect the resonance-like features in Fig. 15.
@@ VIII.2 Missing Hadronic Energy
We explore the impact of missing hadronic energy through constructing the alternative response matrix with an upper limit on this quantity, starting with only events in which all hadronic energy is visible. The results of this closure test are presented in Fig. 17. The methods displaying the most resilience are the single proton version of the CCQE-like estimator and the SF method, followed by the $W$-based and proton-based estimators. The calorimetric and CCQE-like methods are more sensitive to unobserved hadronic components.
Figure 16: Change in overall bias in neutrino energy as a function of the cut on missing hadronic energy. Methods that incorporate transverse kinematic balance (SF, $W$-based) remain stable, whereas calorimetric and muon-kinematic methods degrade more strongly.
@@
(d) $\Delta m^{2}_{23}$.
Figure 17: Oscillation fits under restrictions on missing hadronic energy. The SF method yields the most consistent $\delta_{CP}$ and $\Delta m^{2}_{23}$ extraction, though with reduced statistics.
@@ VIII.3 Angular Dependence
We also construct the alternative response matrix by cutting on the lepton scattering angle, starting with only including forward going leptons before loosening the cut to include larger angles, corresponding to more inelastic interactions. Fits of the oscillated spectra are shown in Fig. 19, revealing that the $W^{2}$-based estimator is comparatively stable, while the purely CCQE-like approach is highly susceptible to angle mismodelling. The SF method offers the second best performance, likely because this method is applied to selected events that still resemble quasielastic interactions even at large scattering angles.
Figure 18: Change in overall bias in neutrino energy as a function of a cut on lepton scattering angle, comparing all estimators. Forward-going selections reduce the impact of angular mismodelling, with the $W$-based method showing the most stable behavior.
@@
(d) $\Delta m^{2}_{23}$.
Figure 19: Extracted oscillation parameters when restricting events by lepton scattering angle. The $W$-based estimator converges rapidly to the correct $\delta_{CP}$ and $\Delta m^{2}_{23}$ values, while the muon-only approach shows strong dependence on $\theta$.
@@ IX Summary
The $W^{2}$ method consistently exhibits the smallest overall bias as functions of the true energy, with a variance comparable to other inclusive estimators, particularly once resolution effects are considered. Additionally, the $W^{2}$ method shows the weakest dependence on FSI effects, most notably when using the GENIE neutrino event generator. When probing the stability of each estimator against the visible hadronic invariant mass, missing hadronic energy, and the angle and momentum of the scattered lepton, we find that the $W^{2}$ method often shows the least variation in bias with respect to these variables, especially in the case of the lepton’s scattering angle.
@@
In this work, we examine each method individually, highlighting the contexts and underlying motivations for which each estimator performs best. This thorough examination provides a detailed guidance for future experiments on how to combine estimators to enhance overall performance. Certain considerations when combining estimators have not explored in this work, for example migration of events between channels due to reconstruction failures possibly leading to a decline in performance.
In this analysis we calculate the estimated energy event by event using each method assuming the detection thresholds established by existing LArTPCs coupled with perfect efficiency and particle separation performance. Future studies of this subject should attempt to use more complete detector response and reconstruction models to further explore the interplay between the cross section modeling effects explored here and these other aspects of oscillation measurements.